OpenStudy (kaleidoscopicsink):

The position of a particle on the x-axis at time t, t > 0, is s(t) = ln(t) with t measured in seconds and s(t) measured in feet. What is the average velocity of the particle for 1 ≤ t ≤ e? Can someone show me how to do this?

OpenStudy (kaleidoscopicsink):

@Loser66

OpenStudy (kaleidoscopicsink):

@mathmale

OpenStudy (kaleidoscopicsink):

@pooja195

OpenStudy (aihberkhan):

Okay. Looking at the average velocity between \(t=2e\) and \(t=e\). You can see it here: \[ \frac{ \ln(2e) - \ln (2e) }{ 2e - e }\]

OpenStudy (kaleidoscopicsink):

1 e e − 1 1/e-1 I think it's d or b. But I'm not sure.

OpenStudy (aihberkhan):

Well... lets finish the work first. Now let's simplify. First let's simplify the denominator. \(2e - e\) will become \(e(2 - 1)\). Then that will become, simply, \(e\). So we will have: \[\frac{ \ln(2e) - \ln(2e) }{ e }\]

OpenStudy (aihberkhan):

Lastly, you have to us the log property for simplifying the numerator: \[\ln(a) -\ln (b) = \ln (\frac{ a }{ b })\]

OpenStudy (aihberkhan):

Try solving.

OpenStudy (kaleidoscopicsink):

1 e e − 1 1/e-1 These are the answer choices. Give me a second to attempt solving.

OpenStudy (aihberkhan):

Okay take your time! :)

OpenStudy (kaleidoscopicsink):

if you had ln(a/b) that would go to ln(2e/2e) so ln 1 which is 1/x therefore being d?

OpenStudy (aihberkhan):

Great Job! :) I would say that as well! :)

OpenStudy (aihberkhan):

Hope this helped! Have a great day! :) A medal would also be appreciated! Just click best response next to my name! Also, a fan would be appreciated as well! Just hover over my icon and click "Become A Fan". This will allow you to see every time I am online! :) If you see that I am online and need help with a question, just tag me in your question! @KaleidoscopicsInk

OpenStudy (kaleidoscopicsink):

Thank you so much for your help! I fanned you and will take you up on that offer most likely. cx Have a great day.